Problem 17-32 (LO. 3, 7) Benton Company (BC), a calendar year entity, has one owner, who is in the 37% Federal income tax bracket (any net capital gains or dividends would be taxed at a 20% rate). BC's gross income is $395,000, and its ordinary trade or business deductions are $245,000. Ignore the standard deduction (or itemized deductions) and the deduction for qualified business income.

If required, round computations to the nearest dollar.

a. BC is operated as a proprietorship, and the owner withdraws $100,000 for personal use. BC's taxable income for the current year is $150,000, and the tax liability associated with the income from the sole proprietorship is $55,500

b. BC is operated as a C corporation, pays out $100,000 as salary but no dividends to its shareholder. BC's taxable income for the current year is $50,000, and BC's tax liability is $10,500. The shareholder's tax liability is $37,000.

c. BC is operated as a C corporation and pays out no salary or dividends to its shareholder. BC's taxable income for the current year is $150,000, and BC's tax liability is $31,500.

d. BC is operated as a C corporation, pays out $100,000 as salary to its shareholder, and pays out the remainder of its earnings as dividends to its shareholder. BC's taxable income for the current year is $50,000, and BC's tax liability is $10,500. The shareholder's tax liability is $_______.
Answer #1

The questions a, b, c and d are already answered (or there are no questions there). Please find below answer to question d

Answer d:

BC operates as C corporation  pays out the remainder of its earnings as dividends to its shareholder.

So dividend to shareholder = 50000 - 10500 = $39,500

Shareholder earnings:

Salary income = $100,000

Dividend income = $39,500

Shareholder's tax liability = 100000 * 37% + 39500 * 20% = $44,900

The shareholder's tax liability is $44,900